OpenStudy (anonymous):

Can anyone please help me with a math question about asymptotes? I know what to do.. but can you please check? -----> f(x) =x^2-25/x+4 Determine all asymptotes of this function including horizontal, vertical, and oblique (slant) THANKS !!

OpenStudy (nathan917):

@Megan1101 Go here http://www.wolframalpha.com/input/?i=%28x^2-25%29%2F%28x%2B4%29+retricemptotes

OpenStudy (nathan917):

Or. To find the vertical asymptotes, make the denominator zero: x + 4 = 0 x = -4 For the others, consider what happens as x approaches infinity. In that case, the -25 and +4 become insignificant as compared to the x^2 and the x, so we're left with: f(x) = (x^2) / x = x So the slant asymptote is y = x

OpenStudy (anonymous):

Here is my answer: Vertical Asymptotes: x = -4 Horizontal Asymptotes: There is no horizontal Asymptotes. But for Oblique Asymptotes: HOW DO I DIVIDE?

OpenStudy (nathan917):

I'm Not sure. @AriPotta Can you help with some math please?

OpenStudy (anonymous):

ok well how did you get the other answer: y = x?

OpenStudy (jdoe0001):

@Megan1101 since the dividend is of the form (x-a), just use a synthetic division to get the quotient and the slant asymptote

OpenStudy (anonymous):

umm ok well I got 2 different answers! Can you please help? @jdoe0001

OpenStudy (jdoe0001):

$$ ^\boxed{-4} \pmatrix{1 & 0 & -25\\ & -4 & 16\\ --& -- & --\\ 1 & -4 & -9 }\\ \text{which gives me a result of } \boldsymbol{x-4}-\cfrac{9}{x+4}\\ $$ bolded non-rational part of the result is the slant asymptote

OpenStudy (anonymous):

sorry for replying late, thank you! ;)

OpenStudy (anonymous):

Ok I totally understand it now! thanks again!

OpenStudy (jdoe0001):

yw